OpenStudy (anonymous):

Find the sum to infinity of the geometric progression with the first 3 terms 0.5, 0.125, 0.03125

OpenStudy (anonymous):

can you tell what is the common ratio r here?

OpenStudy (anonymous):

Hint: Common ratio r is the ratio of two successive terms in GP.. \[r=\frac{ x _{n+1} }{ x }\]

OpenStudy (anonymous):

common ratio = 0.25..?

OpenStudy (wolf1728):

It seems to me that the series is based on this formula: 1/(2^n) So the first 3 numbers are: 1/2^1 1/2^3 1/2^5

OpenStudy (anonymous):

\[S=\lim_{n \rightarrow \infty}\frac{ a_1-a_{n+1} }{ 1-r }\]

OpenStudy (anonymous):

\[a_1=0.5\]\[\lim_{n \rightarrow \infty}a_{n+1}=0\]\[r=0.25\] and\[S=\frac{ a_1 }{ 1-r}\]

OpenStudy (anonymous):

Simply use the formula that CarlosGP mentioned i.e S= a/ (1-r) where a= first term | r= common ratio S= 0.5/ (1-0.25) = ?